OpenStudy (anonymous):

Using a T! 89 - point slope form if m = -2 and passes through (3,4) find the equation of the line.

jimthompson5910 (jim_thompson5910):

Point slope form \[\Large y - y_1 = m(x-x_1)\] Make the proper substitutions to get \[\Large y - 4 = -2(x-3)\] now solve for y

jimthompson5910 (jim_thompson5910):

now type in `solve(y-4=-2*(x-3),y)` to have the TI-89 solve for y. You'll have to do a slight rearrangement to get it into y = mx+b form

jimthompson5910 (jim_thompson5910):

the `solve(` function is found by hitting F2 in the home screen (to get the "algebra" menu)

OpenStudy (anonymous):

when I try that it gives me an err: dimension.

jimthompson5910 (jim_thompson5910):

because you made x and y be defined based on those matrices we were working with earlier (which is my fault)

jimthompson5910 (jim_thompson5910):

what you have to do is clear the RAM memory to free up x and y

jimthompson5910 (jim_thompson5910):

2nd key ---> 6 to bring up the MEM screen hit F1 and reset the RAM only (all RAM) do not touch the ROM at all because that's where the APPs are stored